Give the inverse, converse and contrapositive for each of the following propositions.

a) If x + y = 1 then x2 + y2 ≥ 1

b) If 2 + 2 = 4 then 3 + 3 = 8


1
Expert's answer
2022-03-25T06:23:28-0400

a) inverse: If "x+y\not=1" , then "x^2+y^2<1"

converse: if "x^2+y^2\u22651" , then x + y = 1

contrapositive: If "x^2+y^2<1" , then "x+y\not=1"


b) inverse: If "2+2\not=4" , then "3+3\not=8"

converse: if "3+3=8" , then 2+2=4

contrapositive: If "3+3\not=8" , then "2+2\not=4"